# Don’t Pay Attention

### v1

@@
Abstract
The Transformer has become the de facto standard for large language models and a wide range of downstream tasks across various domains. Despite its numerous advantages like inherent training parallelism, the Transformer still faces key challenges due to its inability to effectively process sequences beyond a fixed context window and the quadratic complexity of its attention mechanism. These challenges have renewed interest in RNN-like architectures, which offer linear scaling with sequence length and improved handling of long-range dependencies, albeit with limited parallelism due to their inherently recurrent nature. In this paper, we propose Avey, a new neural foundational architecture that breaks away from both attention and recurrence. Avey comprises a ranker and an autoregressive neural processor, which collaboratively identify and contextualize only the most relevant tokens for any given token, regardless of their positions in the sequence. Specifically, Avey decouples sequence length from context width, thus enabling effective processing of arbitrarily long sequences. Experimental results show that Avey compares favorably to the Transformer across a variety of standard short-range NLP benchmarks, while notably excelling at capturing long-range dependencies.
@@ 1 Introduction
The Transformer [98] has emerged as one of the most influential AI innovations in recent years, profoundly impacting various aspects of modern life, including work, science, and art, to mention just a few. Notably, Large Language Models (LLMs) are almost universally based on the Transformer [25], which has demonstrated remarkable performance in natural language processing (NLP) [66, 55, 75] and various other fields [32, 56, 2].
@@
The Transformer’s state-of-the-art performance is largely driven by its recurrence-free self-attention mechanism, which enables parallel processing of entire token sequences. Nevertheless, the computational and memory costs of self-attention grow quadratically with sequence length, making it inefficient for handling arbitrarily long sequences. Extensive research has been conducted over the years to address this limitation [94], with a noticeable emphasis on linearizing attention [48, 9, 106, 99]. These linear approaches aim at approximating self-attention in a more computationally efficient manner, without considerably compromising performance.
@@
Nonetheless, linear attention mechanisms have generally underperformed the original self-attention mechanism, often by a significant margin in language modeling tasks [102, 47]. While recent linear models such as RWKV [70] and RetNet [91] have shown promising results, substantial progress is still needed before they can reliably and consistently surpass the Transformer [52]. In addition, these models have yet to demonstrate definitive empirical effectiveness at scale [25]. This persistent performance gap between quadratic and linear approaches has spurred renewed interest in RNN-based architectures, which offer linear scalability with sequence length but limit parallelism due to their inherently cyclical nature.
@@
To exemplify, state space models (SSMs) [44, 28], which are viewed as extensions of RNNs, have recently emerged as a compelling class of architectures. Unlike traditional RNNs, SSMs can parameterize their state transition matrices in a structured manner (e.g., via using a diagonal plus low-rank decomposition) to improve computational efficiency and enhance gradient flow. A specialized subclass of these models, known as structured state space sequence (S4) models [27, 28], has garnered growing attention. Yet, despite their theoretical appeal, S4 models struggled with language modeling tasks, typically trailing Transformers by several points [20, 27].
@@
Most recently, Mamba [25] advanced S4 models by enhancing their selectivity and effectiveness while enabling high training concurrency. It demonstrated strong performance on tasks involving long-range dependencies and compared favorably to Transformers in language modeling. However, training, scaling, and interpreting Mamba—and SSMs more broadly [87, 72, 31]—remain challenging, while continue to be promising [13].
@@
This paper introduces Avey [^1] Avey is not an acronym, but a name that the authors like., a new architecture for language modeling that entirely departs from Transformer-based and RNN-like designs. Avey is a flexible, sequence-length-invariant model that decouples sequence length from context width, thus enabling effective processing of long-range sequences. It preserves the influence of tokens that appear outside its context window, regardless of their positions in the sequence. This is achieved via a weighted-selective-split interaction mechanism, which systematically skips irrelevant tokens beyond the context window and ensures direct interactions with only relevant ones, retaining their contributions irrespective of sequence length.
Figure 1: Needle-in-a-Haystack test performance comparison between Transformer++, Mamba, RWKV-7, and Avey, all using 1.5B parameters. The x-axis denotes the lengths of haystacks (i.e., documents with distractor texts, varying from 2k to 64k tokens) and the y-axis refers to the position of the needle (i.e., a short sentence) within any of the haystacks. A green cell means that the model was able to recall the needle and a red one entails that it could not. Transformer++, Mamba, and RWKV-7 were all trained with 2k-token contexts, while Avey was trained with only a 512-token context and was able to extrapolate to the maximum needed modeling capacity.
Fig. 1 illustrates Avey’s ability to generalize beyond its training context. A popular benchmark for evaluating this capability is Needle-in-a-Haystack (NiaH) [45]. This benchmark measures a model’s capacity to recite a specific sentence (i.e., the needle) placed at an arbitrary position within a large body of distractor text (i.e., the haystack). Since its introduction, NiaH has become a widely used sandbox for probing the limits of long-context language models in capturing distant dependencies, and smaller models in generalizing beyond their trained context windows [22]. As shown in the figure, Transformer++ (i.e., the Transformer with an enhanced architecture and training recipe– see Section 3.1), which was trained with a 2k-token context window, could not generalize beyond that limit. In contrast, Mamba and RWKV-7 [70], also trained with 2k-token context windows, managed to generalize to nearly 8k and 16k tokens, respectively. Most notably, Avey, despite being trained on a context window of only 512 tokens, successfully generalized to the maximum tested sequence length of 64k tokens, demonstrating strong extrapolative capability far beyond its original training regime.
@@
  • In contrast to the Transformer, we demonstrate that Avey can scale far beyond its context window using the standard Single Needle-In-A-Haystack (S-NIAH) benchmark suite from RULER [38].
@@
  • We provide a comprehensive survey of related work in Appendix O.
@@ 2 Avey
As indicated earlier, Avey comprises two components, a ranker and an autoregressive neural processor. Below, we provide a detailed explanation of each component.
@@ 2.1 Ranker
The current split is defined as the one that either contains the token to be predicted (e.g., split 4 may contain only embedding 7, and Avey will aim to predict token 8) or contributes to predicting the first token in its subsequent split (e.g., split 4 serves in predicting token 9, which will belong to split 5 once predicted). To determine relevance, the ranker computes a similarity score between the current split, say $S_{c}$, and each preceding split, say $S_{p}$, using the MaxSim operator [49], originally proposed and utilized in Information Retrieval. Specifically, pairwise similarities are calculated (e.g., using a cosine function) between each embedding in $S_{c}$ and all embeddings in $S_{p}$. For each embedding in $S_{c}$, the maximum similarity across all $S_{p}$’s embeddings is taken, and then the maxima of all $S_{c}$’s embeddings are added to yield the final MaxSim score (see Fig. 2). This score signifies how relevant $S_{p}$ is to $S_{c}$.
@@
As a result, the weighted-selective-split interaction mechanism does not only allow the ranker to rank splits based on relevance but also the neural processor to contextualize them accordingly, as each selected split is pre-weighted by its relevance score. This empowers the neural processor to judiciously leverage global information (i.e., splits beyond the context width) by focusing selectively on only relevant features, emphasizing informative ones and deemphasizing less useful ones, thus enhancing performance. We analyze the impact of weighting the top k splits in Appendix K.
Note that the ranker is invoked only once per full forward and backward pass. To elucidate, Avey’s depth can be increased by stacking multiple layers within the neural processor (see Fig. 3), thereby enabling the modeling of complex, hierarchical patterns. In contrast, only a single ranker is required before the stack of layers within the processor, regardless of their number. Once the ranker identifies the top k relevant splits of the current split, the neural processor contextualizes them all using one or more layers.
Consequently, during training, each current split is matched once against every preceding split. This results in a compute cost of $\frac{N/S\cdot(N/S+1)}{2}\cdot S^{2}d$ or a time complexity of $O(N^{2}d)$, where $N$ is the sequence length, $S$ is the split size, and $d$ is the embedding dimension. This complexity assumes that scalar multiply-add operations (e.g., those used in computing cosine similarity for MaxSim) and comparisons (e.g., those utilized to determine maximum scores) are constant-time. We next discuss the neural processor.
@@ 2.2.1 The Enricher
The enricher aims at enriching the quality of each token representation via expanding the quantity of its learnable features, thereby enabling the contextualizer to capture more nuanced distinctions between tokens. Concretely, it is a one-layer, position-wise neural network (i.e., the input to each neuron is a single scalar element from an embedding), thus operating on each embedding independently, without considering neighboring embeddings. As such, it allows intra-feature interactions within the context of each individual embedding, facilitating the learning of higher-order and more expressive representations. The enricher can expand each input embedding by an arbitrary factor. We study the effect of varying the expansion factor on Avey’s performance in Appendix D, and ablate the enricher’s contribution in Appendix K.
@@
As demonstrated in Fig. 3, the enricher feeds both, the contextualizer and the fuser. In particular, it bypasses a portion of each expanded embedding directly to the fuser in a technique that we refer to as partial-embedding bypassing. More precisely, the output of the enricher, $\mathbf{Z}\in\mathbb{R}^{C\times m}$, is split into two parts: (1) the head $\mathbf{Z}_{h}\in\mathbb{R}^{C\times m_{h}}$, which is bypassed directly to the fuser, and (2) the tail $\mathbf{Z}_{t}\in\mathbb{R}^{C\times m_{t}}$, which is forwarded to the contextualizer, where $m=m_{h}+m_{t}$. Consequently, varying the tail size alters the head size, which can influence Avey’s performance. We investigate the impact of different tail sizes on Avey’s performance in Appendix E.
The partial-embedding bypassing technique allows preserving raw distinctive features of each embedding, thus inducing representations with inherent diversity. This diversity may serve in alleviating issues like entropy collapse [107], where the contextualizer increasingly focuses on a few tokens, and over-smoothing [109, 85, 110], where embeddings become increasingly similar, as Avey’s depth is increased. We analyze the significance of partial-embedding bypassing on Avey’s effectiveness in Appendix K.
Lastly, Equation 1 implies that each neuron performs a weighted sum of $d$ input features (i.e., the elements of an embedding), incurring $d-1$ multiply-add operations. Since $d$ is projected to a higher dimension $m$ [^3] In our case, we experiment with $m$ being a multiple of $d$, entailing that $m\geq 2d$ (see Appendix D)., the total computational cost is $m(d-1)$ per token. For a sequence of $N$ tokens, the cost becomes $Nm(d-1)$, or asymptotically $\mathcal{O}(Nmd)$.
@@ 2.2.2 The Contextualizer
The $m_{t}$ part of each enriched embedding is further divided into two equal portions, $m_{tl}$ (or left portion) and $m_{tr}$ (or right portion), to enable judicious control of information flow through the neural processor. Specifically, $m_{tl}$ serves as a gating mechanism for $m_{tr}$, regulating how much of its contextualized feature values are propagated forward. Both $m_{tl}$ and $m_{tr}$ are learnable by the model, hence, allowing $m_{tl}$ to dynamically capture the significance of each $m_{tr}$’s feature, and emphasize or deemphasize its influence accordingly. This gating mechanism was inspired from gMLP [53] and resembles that of Gated Linear Units [15, 84, 100].
@@
Equation 2 suggests that each neuron in the contextualizer’s network performs a weighted sum of the cosine similarities between embeddings (denoted by $\mathcal{N}(\mathbf{Z}_{tr})\mathcal{N}(\mathbf{Z}_{tr}^{\top})$) and the embeddings themselves (denoted by $\mathbf{Z}_{tr}$). This introduces a level of selectivity into the neural processor, as described and advocated by [25]. Specifically, it makes the parametrization of the neural processor dynamic, enabling it to disregard or focus on information during inference based on the input. We examine the influence of dynamic parametrization on Avey’s performance in Appendix K.
@@ 2.2.3 The Fuser
The fuser is designed to learn an optimal function, referred to as fusion [^4] The name is inspired from the CNN literature [40]., that integrates uncontextualized features (i.e., those of dimension $m_{h}$, bypassed by the partial-embedding bypassing technique presented in Section 2.2.1) with contextualized features (i.e., those of dimension $m_{t}/2$, output by the contextualizer). Subsequently, it produces, for each input token, a contracted representation that matches the token’s original embedding dimension $d$ (see Fig. 3). Akin to the enricher, it is a one-layer, position-wise neural network, which operates on each embedding of dimension $m_{h}+m_{t}/2$ independently.
[…]
@@
Considering the aggregate computational costs of the ranker, enricher, contextualizer, and fuser, Avey exhibits a training time complexity of $\mathcal{O}(L(2Nmd+NkCm_{t})+N^{2}d)$, where $L$ denotes the number of neural processor layers. As the term $N^{2}d$ dominates asymptotically, the overall complexity simplifies to $\mathcal{O}(N^{2}d)$. During inference, the complexity reduces to $\mathcal{O}(N)$, or linear per token. We elaborate on Avey’s inference time complexity in Appendix N.
Figure 4: The Time to First Token (TTFT) for Avey, Transformer++, Mamba, and RWKV-7 using $\sim$1.5B parameters, across varying sequence lengths. See Section 3.1 for details on the experimental methodology.
To characterize Avey’s inference-time efficiency relative to other models, we benchmarked Time to First Token (TTFT)—a key latency metric for real-time applications [36, 54, 17]—on a single NVIDIA H200 GPU. The evaluation included Avey, Transformer++, Mamba, and RWKV-7 across a range of sequence lengths. As shown in Figure 4, Transformer++ demonstrates an approximately quadratic increase in TTFT as the sequence length $N$ grows, owing to its full self-attention mechanism, which must operate over the entire prompt before generating the first token. In contrast, Mamba and RWKV-7 scale linearly with $N$, as they require a full forward pass to construct their RNN-style hidden states prior to emitting the first token. While Avey is also theoretically expected to scale linearly, its empirical TTFT is significantly lower than that of Transformer++, Mamba, and RWKV-7. This efficiency gain stems from Avey’s architectural design, which invokes its dominant contributor to inference complexity (see Appendix N), namely, the ranker, only once per forward pass. As a result, the ranker imposes minimal computational overhead in practice, enabling Avey to achieve substantially lower TTFT and making it particularly well-suited for latency-sensitive, real-world applications (e.g., conversational AI and edge deployments).

### v2

@@
Don’t Pay Attention
[…]
svg-inkscape/.
Don’t Pay Attention Mohammad Hammoud,  Avey AI,  mhh@avey.ai Devang Acharya,  Avey AI,  dacharya@avey.ai The Transformer has become the de facto standard for modern language models owing to its parallelizable training and effective autoregressive decoding. However, its fixed context window and the quadratic time and memory costs of its self-attention mechanism remain central bottlenecks. These constraints have revived interest in recurrent architectures that scale linearly with sequence length, but at the cost of reduced parallelism. In this paper, we introduce Avey, a new foundational architecture that breaks away from both attention and recurrence. Avey pairs a ranker with an autoregressive neural processor to select and contextualize only the most relevant tokens for any given token. Specifically, it decouples sequence length from context width, thus enabling effective and efficient processing of arbitrarily long sequences. Results show that Avey compares favorably to the Transformer across a variety of standard short-range NLP benchmarks, while significantly outperforming it on tasks requiring long-range dependency modeling. Date: November 14, 2025 Website: https://avey.ai/research Equal contribution.
@@ 1 Introduction
The Transformer (vaswani2017attention) has emerged as one of the most influential AI innovations in recent years, profoundly impacting various aspects of modern life, including work, science, and art, to mention just a few. Notably, Large Language Models (LLMs) are almost universally based on the Transformer (gu2023mamba), which has demonstrated remarkable performance in natural language processing (NLP) (ouyang2022training; liu2019roberta; raffel2020exploring) and various other fields (he2022masked; liu2021swin; baevski2020wav2vec).
@@
The Transformer’s state-of-the-art performance is largely driven by its recurrence-free self-attention mechanism, which enables parallel processing of entire token sequences. Nevertheless, the computational and memory costs of self-attention grow quadratically with sequence length, making it inefficient for handling arbitrarily long sequences. Extensive research has been conducted over the years to address this limitation (tay2022efficient), with a noticeable emphasis on linearizing attention (katharopoulos2020transformers; choromanski2020rethinking; zhai2021attention; wang2020linformer). These linear approaches aim at approximating self-attention in a more computationally efficient manner, without considerably compromising performance.
@@
Nonetheless, linear attention mechanisms have generally underperformed the original self-attention mechanism, often by a significant margin in language modeling tasks (yang2023gated; kasai2021finetuning). While recent linear models such as RWKV (peng2025rwkv) and RetNet (sun2023retentive) have shown promising results, substantial progress is still needed before they can reliably and consistently surpass the Transformer (li2024survey). In addition, these models have yet to demonstrate definitive empirical effectiveness at scale (gu2023mamba). This persistent performance gap between quadratic and linear approaches has spurred renewed interest in RNN-based architectures, which offer linear scalability with sequence length but limit parallelism due to their inherently cyclical nature.
@@
To exemplify, state space models (SSMs) (kalman1960new; gu2021combining), which are viewed as extensions of RNNs, have recently emerged as a compelling class of architectures. Unlike traditional RNNs, SSMs can parameterize their state transition matrices in a structured manner (e.g., via using a diagonal plus low-rank decomposition) to improve computational efficiency and enhance gradient flow. A specialized subclass of these models, known as structured state space sequence (S4) models (gu2021efficiently; gu2021combining), has garnered growing attention. Yet, despite their theoretical appeal, S4 models struggled with language modeling tasks, typically trailing Transformers by several points (fu2022hungry; gu2021efficiently).
@@
Most recently, Mamba (gu2023mamba) advanced S4 models by enhancing their selectivity and effectiveness while enabling high training concurrency. It demonstrated strong performance on tasks involving long-range dependencies and compared favorably to Transformers in language modeling. However, training, scaling, and interpreting Mamba—and SSMs more broadly (smith2022simplified; poli2023hyena; hasani2022liquid)—remain challenging, while continue to be promising (dao2024transformers).
@@
This paper introduces Avey [^1] Avey is not an acronym, but a name that the authors like., a new architecture for language modeling that departs from Transformer-based and RNN-like designs. Avey is a flexible, sequence-length-invariant model that decouples sequence length from context width, thus enabling effective processing of long-range sequences. It preserves the influence of tokens that appear outside its context window, regardless of their positions in the sequence. This is achieved via a weighted-selective-split interaction mechanism, which systematically skips irrelevant tokens beyond the context window and ensures direct interactions with only relevant ones, retaining their contributions irrespective of sequence length.
Figure 1: Needle-in-a-Haystack test performance comparison between Transformer++, Mamba, RWKV-7, and Avey, all using 1.5B parameters. The x-axis denotes the lengths of haystacks (i.e., documents with distractor texts, varying from 2k to 64k tokens) and the y-axis refers to the position of the needle (i.e., a short sentence) within each of the haystacks. A green cell indicates successful needle recall, while a red cell indicates failure. Transformer++, Mamba, and RWKV-7 were trained with 2k-token context windows, whereas Avey was trained with only a 512-token window yet was able to extrapolate to the longest sequences evaluated.
Fig. 1 illustrates Avey’s ability to generalize beyond its training context. A popular benchmark for evaluating this capability is Needle-in-a-Haystack (NiaH) (kamradt2023needle). This benchmark measures a model’s capacity to recite a specific sentence (i.e., the needle) placed at an arbitrary position within a large body of distractor text (i.e., the haystack). Since its introduction, NiaH has become a widely used sandbox for probing the limits of long-context language models in capturing distant dependencies, and smaller models in generalizing beyond their trained context windows (fu2024data). As shown in the figure, Transformer++ (i.e., the Transformer with an enhanced architecture and training recipe– see Section 3.1 for details), which was trained with a 2k-token context window, could not generalize beyond that limit. In contrast, Mamba and RWKV-7 (peng2025rwkv), also trained with 2k-token windows, managed to generalize to nearly 8k and 16k tokens, respectively. Most notably, Avey, despite being trained on a context window of only 512 tokens, successfully generalized to the maximum tested sequence length of 64k tokens, demonstrating strong extrapolative capability far beyond its original training regime.
@@
  • In contrast to the Transformer, we demonstrate that Avey can scale far beyond its context window using the standard Single Needle-In-A-Haystack (S-NIAH) benchmark suite from RULER (hsieh2024ruler).
@@ 2 Avey
As indicated earlier, Avey comprises two components, a ranker and a neural processor. We next describe each component in detail (see Appendix G for more design intuitions behind them).
@@ 2.1 Ranker
The current split is defined as the one that either contains the token to be predicted (e.g., split 4 may contain only embedding 7, and Avey will aim to predict token 8) or contributes to predicting the first token in its subsequent split (e.g., split 4 serves in predicting token 9, which will belong to split 5 once predicted). To determine relevance, the ranker computes a similarity score between the current split, say $S_{c}$, and each preceding split, say $S_{p}$, using the MaxSim operator (khattab2020colbert), originally proposed and utilized in Information Retrieval. Specifically, pairwise similarities are calculated (e.g., using a cosine function) between each embedding in $S_{c}$ and all embeddings in $S_{p}$. For each embedding in $S_{c}$, the maximum similarity across all $S_{p}$’s embeddings is taken, and then the maxima of all $S_{c}$’s embeddings are added to yield the final MaxSim score (see Fig. 2). This score signifies how relevant $S_{p}$ is to $S_{c}$.
@@
As a result, the weighted-selective-split interaction mechanism does not only allow the ranker to rank splits based on relevance but also the neural processor to contextualize them accordingly, as each selected split is pre-weighted by its relevance score. This empowers the neural processor to judiciously leverage global information (i.e., splits beyond the context width) by focusing selectively on only relevant features, emphasizing informative ones and deemphasizing less useful ones, thus enhancing performance. We analyze the impact of weighting the top k splits in Section 3.3.
Note that the ranker is invoked only once per full forward and backward passes [^3] It is worth emphasizing that the ranker is an internal module that operates solely over in-sequence splits already present in the input. In other words, it does not access external corpora or indexes, and therefore introduces neither retrieval latency nor corpus-freshness dependencies. Consequently, it is not a RAG component, which retrieves external evidence at inference (and/or training) time from a separate knowledge base. The two approaches are orthogonal in fact (the ranker allocates internal context, while RAG modifies the available evidence set). See Appendix E for a detailed discussion.. To elucidate, Avey’s depth can be increased by stacking multiple layers within the neural processor (see Fig. 3), thereby enabling the modeling of complex, hierarchical patterns. In contrast, only a single ranker is required before the stack of layers within the processor, regardless of their number. Once the ranker identifies the top k relevant splits of the current split, the neural processor contextualizes them all using one or more layers.
Consequently, during training, each current split is matched once against every preceding split. This results in a compute cost of $\frac{N/S\cdot(N/S+1)}{2}\cdot S^{2}d$ or a time complexity of $O(N^{2}d)$, where $N$ is the sequence length, $S$ is the split size, and $d$ is the embedding dimension. This complexity assumes that scalar multiply-add operations (e.g., those used in computing cosine similarity for MaxSim) and comparisons (e.g., those utilized to determine maximum scores) are constant-time.
@@ 2.2.1 The Enricher
The enricher aims at enriching the quality of each token representation via expanding the quantity of its learnable features, thereby enabling the contextualizer to capture more nuanced distinctions between tokens. Concretely, it is a one-layer, position-wise neural network (i.e., the input to each neuron is a single scalar element from an embedding), thus operating on each embedding independently, without considering neighboring embeddings. As such, it allows intra-feature interactions within the context of each individual embedding, facilitating the learning of higher-order and more expressive representations. The enricher can expand each input embedding by an arbitrary factor. We study the effect of varying the expansion factor on Avey’s performance in Section 3.2.3, and ablate the enricher’s contribution in Section 3.3.
@@
As demonstrated in Fig. 3, the enricher feeds both, the contextualizer and the fuser. In particular, it bypasses a portion of each expanded embedding directly to the fuser in a technique that we refer to as partial-embedding bypassing. More formally, the output of the enricher, $\mathbf{Z}\in\mathbb{R}^{C\times m}$, is split into two parts: (1) the head $\mathbf{Z}_{h}\in\mathbb{R}^{C\times m_{h}}$, which is bypassed directly to the fuser, and (2) the tail $\mathbf{Z}_{t}\in\mathbb{R}^{C\times m_{t}}$, which is forwarded to the contextualizer, where $m=m_{h}+m_{t}$. Consequently, varying the tail size alters the head size, which can influence Avey’s performance. We investigate the impact of different tail sizes on Avey’s performance in Section 3.2.4.
The partial-embedding bypassing technique allows preserving raw distinctive features of each embedding, thus inducing representations with inherent diversity. This diversity may serve in alleviating issues like entropy collapse (zhai2023stabilizing), where the contextualizer increasingly focuses on a few tokens, and over-smoothing (zhou2021deepvit; shi2022revisiting; zhou2024value), where embeddings become increasingly similar, as Avey’s depth is increased. We analyze the significance of partial-embedding bypassing on Avey’s effectiveness in Section 3.3.
Lastly, Equation 1 implies that each neuron performs a weighted sum of $d$ input features (i.e., the elements of an embedding), incurring d1d-1 multiply-add operations. Since $d$ is projected to a higher dimension mm [^4] In our case, we experiment with $m$ being a multiple of $d$, entailing that $m\geq 2d$ (see Section 3.2.3)., the total computational cost is m(d1)m(d-1) per token. For a sequence of $N$ tokens, the cost becomes Nm(d1)Nm(d-1), or asymptotically $\mathcal{O}(Nmd)$.
@@ 2.2.2 The Contextualizer
The $m_{t}$ part of each enriched embedding is further divided into two equal portions, $m_{tl}$ (or left portion) and $m_{tr}$ (or right portion), to enable judicious control of information flow through the neural processor. Specifically, $m_{tl}$ serves as a gating mechanism for $m_{tr}$, regulating how much of its contextualized feature values are propagated forward. Both $m_{tl}$ and $m_{tr}$ are learnable by the model, hence, allowing $m_{tl}$ to dynamically capture the significance of each $m_{tr}$’s feature, and emphasize or deemphasize its influence accordingly. This gating mechanism was inspired from gMLP (liu2021pay) and resembles that of Gated Linear Units (dauphin2017language; shazeer2020glu; wu2019pay).
@@
Equation 2 suggests that each neuron in the contextualizer’s network performs a weighted sum of the cosine similarities between embeddings (denoted by $\mathcal{N}(\mathbf{Z}_{tr})\mathcal{N}(\mathbf{Z}_{tr}^{\top})$) and the embeddings themselves (denoted by $\mathbf{Z}_{tr}$). This introduces a level of selectivity into the neural processor, as advocated by (gu2023mamba). Specifically, it makes the parametrization of the neural processor dynamic, enabling it to disregard or focus on information during inference based on the input. We examine the influence of dynamic parametrization on Avey’s performance in Section 3.3555See also a discussion on neural contextualization versus attention in Appendix F..
@@ 2.2.3 The Fuser
The fuser is designed to learn an optimal function, referred to as fusion [^6] The name is inspired from the CNN literature (hu2018squeeze)., that integrates uncontextualized features (i.e., those of dimension $m_{h}$, bypassed by the partial-embedding bypassing technique) with contextualized features (i.e., those of dimension $m_{t}/2$, output by the contextualizer). Subsequently, it produces, for each input token, a contracted representation that matches the token’s original embedding dimension $d$ (see Fig. 3). Akin to the enricher, it is a one-layer, position-wise neural network, which operates on each embedding of dimension $m_{h}+m_{t}/2$ independently.
Equation 3 provides a mathematical definition of the fuser, where $\mathbf{Z}_{h}\in\mathbb{R}^{C\times m_{h}}$ (as described in Section 2.2.1) and $\mathbf{c}(\mathbf{Z}_{t})\in\mathbb{R}^{C\times(m_{t}/2)}$ (as suggested by Equation 2) are concatenated, and $\mathbf{O}\in\mathbb{R}^{(m_{h}+m_{t}/2)\times d}$ is a learnable weight matrix representing a linear projection from dimension $m_{h}+m_{t}/2$ back to dimension $d$, where $d<m_{h}+m_{t}/2$ [^7] This inequality will always hold if $m_{h}+m_{t}\geq 2d$, as is the case in our experiments (see Section 3)..
@@
Considering the aggregate computational costs of the ranker, enricher, contextualizer, and fuser, Avey exhibits a training time complexity of $\mathcal{O}(L(2Nmd+NkCm_{t})+N^{2}d)$, where $L$ denotes the number of neural processor layers. As the term $N^{2}d$ dominates asymptotically, the overall complexity simplifies to $\mathcal{O}(N^{2}d)$. During inference, the complexity reduces to $\mathcal{O}(N)$, or linear per token. We provide a detailed analysis of Avey’s time complexity in Appendix D.
Figure 4: The Time to First Token (TTFT) for Avey, Transformer++, Mamba, and RWKV-7 across varying sequence lengths.
Lastly, to characterize Avey’s inference-time efficiency relative to other models, we benchmarked Time to First Token (TTFT)—a key latency metric for real-time applications (horton2024kv; liu2025speculative; dexter2025llm)—on a single NVIDIA H200 GPU. The evaluation included Avey, Transformer++, Mamba, and RWKV-7 across a range of sequence lengths. As shown in Fig. 4, Transformer++ demonstrates an approximately quadratic increase in TTFT as the sequence length $N$ grows, owing to its full self-attention mechanism, which must operate over the entire prompt before generating the first token. In contrast, Mamba and RWKV-7 scale linearly with $N$, as they require a full forward pass to construct their RNN-style hidden states prior to emitting the first token. While Avey is also theoretically expected to scale linearly, its empirical TTFT is significantly lower than that of Transformer++, Mamba, and RWKV-7. This stems from Avey’s architectural design, which invokes its dominant contributor to inference complexity (see Appendix D), namely, the ranker, only once per forward pass. As a result, the ranker imposes minimal computational overhead in practice, enabling Avey to achieve substantially lower TTFT and making it particularly well-suited for latency-sensitive, real-world applications (e.g., conversational AI and edge deployments).